OpenStudy (anonymous):

Solve the following equation, answer as a reduced, mixed number. 4(7x - 1) = 3(2x + 1) - 4(x + 6)

OpenStudy (anonymous):

28x-4=6x+3-4x-24 26x=-17 x=-17/26

OpenStudy (anonymous):

Thanks! Want to help with a couple of others I'm stuck on?